Pam is 1/2 as old as her brother Bill. 8 years ago Bill was 10 times older than Pam. How old are they now?
Bill = x
Pam = x/2
Bill = x-8
Pam = x/2-8

Answer by nerdybill(5404)


Pam is 1/2 as old as her brother Bill. 8 years ago Bill was 10 times older than Pam. How old are they now?
.
Let x = Bill's age
then
x/2 = Pam's age
.
From:"8 years ago Bill was 10 times older than Pam"
x-8 = 10(x/2 - 8)
Multiplying both sides by 2:
2x-16 = 10(x - 16)
2x-16 = 10x - 160
-16 = 8x - 160
144 = 8x
18 years = x (Bill's age)
.
Pam's age:
x/2 = 18/2 = 9 years